Please help me with this problem.
Find a quadratic function with solutions -5+2i and -5-2i and a leading coefficient of 1.

Answer by scott8148(6628)

if s is a solution (root) of a quadratic equation, then (x-s) is a factor

(x-(-5+2i))(x-(-5-2i))=0 ... (x+5-2i)(x+5+2i)=0 ... x^2+5x+2ix+5x+25+10i-2ix-10i-4i^2=0 ... x^2+10x+29=0